Please help me solve this problem:
50,000 -10,000(P) = 25,000(P)
Thanks

Answer by stanbon(57366)

50,000 -10,000(P) = 25,000(P)
50,000 = 35000p
p = 50/35
p = 10/7
================
Cheers,
stan H.